eviant:

PLS HELP!!! A line passes through (–7, –5) and (–5, 4). a. Write an equation for the line in point-slope form. b. Rewrite the equation in standard form using integers.

ThisGirlPretty:

@Zepdrix Can you help ^_^

563blackghost:

Since you are given two points you can find the slope between the two. \(\Large\bf{slope = \frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{4-(-5)}{-5-(-7)} \rightarrow \frac{9}{2}}\) Once this is found you are now able to plug into point slope. \(\large\bf{y-y_{1}=m(x-x_{1})}\) \(\bf{(x_{1},y_{1}) \rightarrow first~point}\) Remember that \(\bf{m}\) equals \(\bf{slope}\). What is the equation?

eviant:

so, y+5=9/2(x+7) for the point-slope form?

Vocaloid:

^ well done now, for part b, you would distribute the 9/2 and then re-write the equation into Ax + By = C form